A Japanese company has a bond outstanding that sells for 91 percent of its ¥100,000 par value. The bond has a coupon rate of 5 percent paid annually and matures in 12 years. 


What is the yield to maturity of this bond? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g, 32.16.) 
Answer #1
Face value ¥                   100,000.00
Current price 91% ¥   91,000.00
Annual coupon payment 5% ¥     5,000.00
Redemption price 100% ¥100,000.00
6.00% 7.00%
Year Cash Flow PV factor = 1/ (1+r)^t PV PV factor = 1/ (1+r)^t PV
0 ¥                    -91,000.00 1.000 ¥-91,000.00 1.000 ¥-91,000.00
1 ¥                        5,000.00 0.943 ¥    4,716.98 0.935 ¥    4,672.90
2 ¥                        5,000.00 0.890 ¥    4,449.98 0.873 ¥    4,367.19
3 ¥                        5,000.00 0.840 ¥    4,198.10 0.816 ¥    4,081.49
4 ¥                        5,000.00 0.792 ¥    3,960.47 0.763 ¥    3,814.48
5 ¥                        5,000.00 0.747 ¥    3,736.29 0.713 ¥    3,564.93
6 ¥                        5,000.00 0.705 ¥    3,524.80 0.666 ¥    3,331.71
7 ¥                        5,000.00 0.665 ¥    3,325.29 0.623 ¥    3,113.75
8 ¥                        5,000.00 0.627 ¥    3,137.06 0.582 ¥    2,910.05
9 ¥                        5,000.00 0.592 ¥    2,959.49 0.544 ¥    2,719.67
10 ¥                        5,000.00 0.558 ¥    2,791.97 0.508 ¥    2,541.75
11 ¥                        5,000.00 0.527 ¥    2,633.94 0.475 ¥    2,375.46
12 ¥                        5,000.00 0.497 ¥    2,484.85 0.444 ¥    2,220.06
12 ¥                   100,000.00 0.497 ¥ 49,696.94 0.444 ¥ 44,401.20
¥       616.16 ¥   -6,885.37
NPV @ 0.06 ¥                           616.16
NPV @ 0.07 ¥                      -6,885.37
Difference in both ¥                        7,501.53
YTM =Lower rate + Difference in rates*(NPV at lower rate)/(Lower rate NPV-Higher rate NPV)
'=6%+ (7%-6%)*(616.15/(616.15+6885.3)
YTM 6.08%